OpenStudy (legomyego180):

Rephrase of my previous question regarding line integrals:

OpenStudy (legomyego180):

Integrate h(x, y) = yi + xj over the circle of radius 5 centered at the origin traversed clockwise.

OpenStudy (legomyego180):

Ive worked out: f(x,y)=(5sin(t), 5cos(t)) f'(x,y)=(5cos(t), -5sin(t)) Do I now dot these?

OpenStudy (legomyego180):

Im confused because in the formula it says something like you parameterize f(x,y) then plug that in for your formula. So you'd have: \(\int\limits_{0}^{1}f(x(t), y(t))ds\) Where ds = \(\sqrt{(\frac{ dx }{ dt })^2+(\frac{ dy }{ dt })^2}\) Im confused because everyone keeps telling me to dot the function with its derivative and I dont really see how that has anything to do with this formula...

OpenStudy (legomyego180):

I forgot to add above: "integrating with respect to t"

OpenStudy (legomyego180):

I realize the answer is 0 as well, but I dont understand the process

OpenStudy (holsteremission):

The line integral is \[\int_C \mathbf F\cdot \mathrm d\vec{r}=\int_Ch(x(t),y(t))\cdot\frac{\mathrm df(x(t),y(t))}{\mathrm dt}\,\mathrm dt\]That is to say, you're "dotting" \(h\) with the derivative of \(f\).

OpenStudy (legomyego180):

So for: \[h(x,y)=xy^2i+2j\] \[r(u)=e^ui+e^{-u}j\] \[u \epsilon [0,1]\] \[\int_C \mathbf F\cdot \mathrm d\vec{r}=\int_Ch(x(t),y(t))\cdot\frac{\mathrm df(x(t),y(t))}{\mathrm dt}\,\mathrm dt\] \[(e^ue^{-2u}, 2) \cdot (?)\] Heres where its throwing me off a bit, im not sure how to do the derivatives.

OpenStudy (legomyego180):

Do I take the derivative of h(x,y) then substitute my t parameters in and multiply by the derivative of the t parameters I plugged in?

OpenStudy (holsteremission):

This looks like a completely different problem. Let's just stick to the first for a moment. You have a vector field \(h(x,y)=y\,\vec{i}+x\,\vec{j}\) to be integrated along the circular path of radius \(5\), so a sufficient parameterization might be \(f(t)=\sqrt5\cos t\,\vec{i}+\sqrt5\sin t\,\vec{j}\), with derivative \(\dfrac{\mathrm d}{\mathrm dt}f(t)=-\sqrt5\sin t\,\mathrm dt\,\vec{i}+\sqrt5\cos t\,\mathrm dt\,\vec{j}\). Then you can also write \(h(t)=\sqrt5\sin t\,\vec{i}+\sqrt5\cos t\,\vec{j}\). Now the line integral would be \[\begin{align*} \int_C h(x,y)\cdot\mathrm df&=-\int_0^{2\pi}(\sqrt5\sin t,\sqrt5\cos t)\cdot(-\sqrt5\sin t,\sqrt5\cos t)\,\mathrm dt\\[1ex] &=-5\int_0^{2\pi}(-\sin^2t+\cos^2t)\,\mathrm dt\\[1ex] &=-5\int_0^{2\pi}\cos2t\,\mathrm dt\\[1ex] &=-\frac{5}{2}\int_0^\pi \cos t\,\mathrm dt=\color{red}0 \end{align*}\]Note the negative sign, since \(C\) is traversed in the clockwise direction, though the ultimate result is independent of that. Does all this make sense?

OpenStudy (holsteremission):

(Actually, ignore the \(\mathrm dt\)s in the derivative of \(f\), they shouldn't be there.)

OpenStudy (holsteremission):

Are you using the same path for the second problem?

OpenStudy (legomyego180):

"This looks like a completely different problem. Let's just stick to the first for a moment. You have a vector field \(h(x,y)=y\,\vec{i}+x\,\vec{j}\) to be integrated along the circular path of radius \(5\), so a sufficient parameterization might be \(f(t)=\sqrt5\cos t\,\vec{i}+\sqrt5\sin t\,\vec{j}\), with derivative \(\dfrac{\mathrm d}{\mathrm dt}f(t)=-\sqrt5\sin t\,\mathrm dt\,\vec{i}+\sqrt5\cos t\,\mathrm dt\,\vec{j}\). Then you can also write \(h(t)=\sqrt5\sin t\,\vec{i}+\sqrt5\cos t\,\vec{j}\)." Sure lets stick to the first problem. why is f(t)=\(\sqrt{5}\cos(t)i+\sqrt{5}\sin(t)j\) if the function is \(yi+xj\) and x=rcos(t), and y=rsin(t). Am I thinking about this the wrong way?

OpenStudy (legomyego180):

Sorry about the delayed response

OpenStudy (518nad):

theres no root 5

OpenStudy (518nad):

one possible paramet for circle rad 5 x=5cost y=5sint -pi<t<pi

OpenStudy (holsteremission):

Oops, that's correct. I somehow misinterpreted the question as saying the path was \(x^2+y^2=5\), not \(x^2+y^2=5^2\). Luckily this doesn't change the outcome, though the coefficient in front of the integral would be \(25\) instead of a \(5\).

OpenStudy (holsteremission):

Basically replace every instance of \(\sqrt5\) with \(5\).

OpenStudy (holsteremission):

Just to address your initial confusion, the reason we're using the formula \(\int_C\mathbf F\cdot\mathrm d\vec{r}\) is because this is the line integral formulation over a vector field \(\mathbf F\). The formula you were intent on using, \(\int_Cf\,\mathrm ds\), only applies for scalar fields.

OpenStudy (holsteremission):

And for your second problem, the derivative you would be using is the one of the vector function \(\vec{r}(u)\). \[\vec{r}(u)=e^u\,\vec{i}+e^{-u}\,\vec{j}\implies\frac{\mathrm d\vec{r}(u)}{\mathrm du}=e^u\,\vec{i}-e^{-u}\,\vec{j}\]Given the vector field \(h(x,y)=xy^2\,\vec{i}+2\,\vec{j}\), the line integral over whatever the path you're using is would be \[\int_Ch\cdot\mathrm d\vec{r}=\int_C(e^ue^{-2u},2)\cdot\color{red}{(e^u,-e^{-u})}\,\mathrm du=\int_C(e^{-u},2)\cdot(e^u,-e^{-u})\,\mathrm du\]

OpenStudy (irishboy123):

 i would recommend that you do your best to see what that vector \(\vec {h}(x, y) = yi + xj\) is actually doing over the path it is following. When does it ever point along the path it is following? Plot it at some key points.... And despite that stuff i posted on your other question, viz that the curl is not infallible, it is always worth checking out as a heads up. \(\partial_x \vec F - \partial_y \vec F = \mathbf{????}\). So you might be able to do the answer in yr head before you even start :-)) In terms of doing the integral, namely the first one ..... \(\vec {h}(x, y) = yi + xj\) So you can write \(\oint \color{red}{\left(\begin{matrix}y \\ x\end{matrix}\right)} \cdot \color{blue}{\left(\begin{matrix}dx \\ dy\end{matrix}\right) }= \oint y ~ dx + \oint x ~ dy\), if you like. Or if we are going to parameterise, we should go polar as it is motion in a circle. To polarise the red vector, we say \(x = \cos t, y = \sin t\), so the red vector is \(\left(\begin{matrix} \sin t \\ \cos t\end{matrix}\right)\) for the blue vector, a circle radius 5, we should say \(x = 5 \cos t, y = 5 \sin t\). This means: \(dx = -5\sin t ~ dt, dy = 5 \cos t ~ dt\), and it becomes: \(\oint \color{red}{\left(\begin{matrix} \sin t \\ \cos t\end{matrix}\right)} \cdot 5 \color{blue}{\left(\begin{matrix} - \sin t \\ cos t\end{matrix}\right) } ~ dt \) \(= 5 \oint -\sin^2 t + \cos^2 t ~ dt \) \(= 5 \oint \cos 2t ~ dt , \qquad t \in [0, -2 \pi ]\) this is where we kick in CW motion, ie in the \(\color{red}{-}2 \pi\) thing \( = \frac{5}{2} \left[\sin 2t \right]_{0}^ {-2 \pi }\) \(= 0\)